OpenStudy (anonymous):

n triangle ABC, angle A is a right angle and angle B = 45 degrees. If AB = 19 feet, what is BC?

OpenStudy (jdoe0001):

 45-45-90 rule -> http://ct12345.weebly.com/uploads/9/6/0/9/9609821/805845_orig.gif

OpenStudy (anonymous):

Thank you

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

AB=AC therefore sqrt of (19^2)+(19^2)= 2((19^2))= (appx) 26.87